Circular No. 7165

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 J2 The Lowell Observatory Near-Earth Object Search (LONEOS) reports

its discovery of a comet. The following CCD positions are available:

1999 UT  α_{2000}  δ_{2000}  m_{1} Observer
May 13.40333 16^{h}58^{m}28.^{s}92 +63^{o}43′15′′.4 16.0 Skiff
13.42119 16 58 27.59 +63 43 15.7  _{”}
13.43905 16 58 26.12 +63 43 15.2  _{”}
13.87581 16 57 50.18 +63 43 06.8 15.5 Tichý
13.87745 16 57 50.08 +63 43 07.3  _{”}
13.87869 16 57 49.95 +63 43 06.9  _{”}

B. A. Skiff (LONEOS). 0.59-m LONEOS Schmidt telescope + R filter. Measurer B. W. Koehn. Well-condensed nucleus in a coma of diameter ∼ 15′′, with a faint tail extending ∼ 40′′ in p.a. 30^{o}.

M. Tichý and Z. Moravec (Kleť). 0.57-m f/5.2 reflector. 10′′ coma.

4U 1630–47 M. L. McCollough, Universities Space Research Association and Mar-

shall Space Flight Center (MSFC); B. A. Harmon, MSFC, NASA; S. Dieters, University of Alabama, Huntsville; and R. Wijnands, Astronomical Institute, University of Amsterdam, communicate: “The CGRO BATSE earth-occultation monitoring (20–100 keV) indicates that the recurrent blackhole x-ray transient 4U 1630–47 (cf. IAUC 6389, 6822) has shown increased activity. It has risen over the last 18 days from 70 mCrab (20–100 keV) on Apr. 20 to 140 mCrab on May 7. It has a hard spectrum (powerlaw index −2.6 ± 0.2) with emission extending past 200 keV. A public target-of-opportunity RXTE PCA observations on May 8.05 UT showed the source to be at 75 mCrab (2–20 keV) with a spectrum described by a power law with an index of −1.6 ± 0.06 and no evidence of a blackbody component. A strong quasiperiodic oscillation was found in the data (2–60 keV) with a centroid frequency of 0.855 ± 0.005 Hz, FWHM of 0.13 ± 0.01, and fractional rms amplitude of 16.3 ± 0.5 percent. This renewed activity of 4U 1630–47 follows a significant outburst in Feb. 1998 (IAUC 6822) and does not fall within the expected ∼ 690-day outburst cycle of this source. With an outburst, radio emission is expected, and observations at radio wavelengths is encouraged.”

1999 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green